find the equation of the normal to the curve y=x^2-4x at the point where x=-3. find the points A and B where the normal cuts the x-axis and y-axis respectively. Hence find the exact length of AB
Answer by Alan3354(69443)

find the equation of the normal to the curve y=x^2-4x at the point where x=-3. find the points A and B where the normal cuts the x-axis and y-axis respectively. Hence find the exact length of AB
--------------
Find the y-value at x = -3
Find the slope at x = -3
The slope of the normal is the negative inverse of the curve's slope.
----
Use y-y1 = m*(m-x1) for the equation where (x1,y1) is the point (-3,y)
=====================
email via the TY note for help, or to check your work.